OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (amistre64):

use the dot product method ....

OpenStudy (amistre64):

\[\cos\alpha=\frac{u\cdot v}{|u||v|}\]